OpenStudy (anonymous):

please help me s^4sqrt(y-5=3

OpenStudy (anonymous):

Do you mean this? \[\huge s^4\sqrt{y-5}=3\]

OpenStudy (anonymous):

^4sqrt(y-5=3

OpenStudy (anonymous):

Please use the equation button.